# Bayesian Inference General Procedures for A Single-subject Test Study

### v2

@@ Abstract
Abnormality detection in the identification of a single-subject which deviates from the majority of the dataset that comes from a control group is a critical problem. A common approach is to assume that the control group can be characterised in terms of standard Normal statistics and the detection of single abnormal subject is in that context. But in many situations the control group can not be described in terms of Gaussian statistics and the use of standard statistics is inappropriate. This paper presents a Bayesian Inference General Procedures for A Single-Subject Test (BIGPAST), designed to mitigate the effects of skewness under the assumption that the dataset of control group comes from the skewed Student’s $t$ distribution. BIGPAST operates under the null hypothesis that the single-subject follows the same distribution as the control group. We assess BIGPAST’s performance against other methods through a series of simulation studies. The results demonstrate that BIGPAST is robust against deviations from normality and outperforms the existing approaches in terms of accuracy. This is because BIGPAST can effectively reduce model misspecification errors under the skewed Student’s $t$ assumption. We apply BIGPAST to a MEG dataset consisting of an individual with mild traumatic brain injury and an age and gender-matched control group, demonstrating its effectiveness in detecting abnormalities in the single-subject.
Keywords— Bayesian inference, Skewed Student’s $t$ distribution, Single-subject test, Control group, Magnetoencephalography (MEG), Jeffery’s prior, and False discovery rate
@@ 1 Introduction
Abnormality detection is a fundamental task in many scientific fields, including medicine, psychology, and neuroscience (Scholz and Stephens, 1987; Crawford and Howell, 1998; Crawford and Garthwaite, 2007). One example of such abnormality detection lies in the detection of traumatic brain injury (TBI), which is a particularly common problem, with over 2.8 million individuals a year in America presenting at an Emergency room (Taylor et al., 2017). We exemplify the use of skewed Student’s $t$ statistics to the study of brain activity as a potential biomarker for the identification of a brain injury. We use magnetoencephalography (MEG) as the technology for measuring brain activity and compare single individual to normal controls from age and gender-matched cohorts. Existing literature normally assumes that the dataset from the control group follows a Normal distribution. However, extensive practical evidence shows that the naive Normal distribution assumption is often violated. This violation is frequently observed in brain Magnetoencephalography (MEG) datasets. Therefore, the current approaches using Normal statistics could not be directly applied. Unlike the existing literature, we assume that the dataset from the control group belongs to a more general distributional family - the skewed Student’s $t$ distribution, which includes normal distribution and Student’s $t$ as a special case. The benefit of this skewed Student’s $t$ distributional assumption is that it could cover a variety of empirical behaviours and is thus more suitable for practical needs. This generalisation is essential since many brain Magnetoencephalography (MEG) datasets show a significant level of asymmetry in the observation distribution and hence are highly skewed. It is also the case that the high skewness not only appears in a MEG dataset, but also in other fields, such as cybersecurity (Kim and Song, 2023), medicine (Lu and Huang, 2017), and neuroscience (Buzsáki and Mizuseki, 2014) to name only a few.
Under the assumption of the skewed Student’s $t$ distribution in this article, we describe the statistical issues involved in the comparison of a single individual with a large normal control group. The statistical issues arise from the interplay between observation from single-subject(whether well-mixed or not), the type of hypothesis (two-sided, less, greater), and the sign of skewness parameter which we will discuss one-by-one in the following paragraphs.
Under the assumption that dataset observed from the control group follows a Normal distribution, there exist several methods that are able to compare a single-subject against a control group. Crawford and Howell (1998) proposed a $t$-score test defined as follows:
@@
where $x^{*}$ represents the observation of the single-subject, $\bar{x}$ denotes the average of $n$ observations from the control group, and $s$ stands for the standard deviation of the dataset from the control group. The test score $t$ follows a Student’s $t$ distribution with $n-1$ degrees of freedom. Let $\alpha$ be the prespecified significance level, define the $p$ value as $\mathbb{P}(t>|t_{1-\alpha/2}(n-1)|)$ where $t_{1-\alpha/2}(n-1)$ is the $(1-\alpha/2)$-th critical value of Student’s $t$ distribution with $n-1$ degrees of freedom. If the $p$ value is less than $\alpha$, then one can claim that $x^{*}$ has an abnormality. Moreover, under the assumption of skewed Student’s $t$ distribution, Crawford et al. (2006) examined the effects of departures from normality on testing for a deficit in a single-subject via using a Leptokurtic distribution with different skewness parameters. They asserted that the $t$ score (Crawford and Howell, 1998) outperforms the $z$ score in terms of Type I error, and the $t$ score shows robustness as good as the result under the Normal assumption. Additionally, Crawford and Garthwaite (2006) investigated the statistical power of testing for an abnormality in a single-subject using Monte Carlo simulations. They claimed that the power of $t$ score performs slightly worse compared with the results under Normal assumption.
The simulation results for Type I error in Crawford et al. (2006) and power in Crawford and Garthwaite (2006) led to the assertion that the $t$ score (Crawford and Howell, 1998) is robust even when the underlying distribution is skewed Student’s $t$ distribution. However, the simulation settings in Crawford et al. (2006)Crawford and Garthwaite (2006) are not comprehensive (ground true observations of a single subject are not well-mixed) to evaluate the performance of $t$ score under skewed Student’s $t$ distribution.
@@
Table 1: The data generation mechanism of single-subject observations in Crawford et al. (2006)Crawford and Garthwaite (2006) respectively. ‘NA’ means that the observations in the cell is not available. ‘Positive’ means that the single-subject is claimed as abnormal, while ‘Negative’ means that the single-subject is claimed as normal.
@@
For instance, in evaluating Type I error, Crawford et al. (2006) generates all single-subject observations from the distribution of the control group, see Table 1(a). In  Table 1(a), $n_{01}$ is the number of false positive (FP) while $n_{00}$ is the number of true negative (TN). The false positive rate (FPR, i.e., Type I error) is defined as $n_{01}/(n_{01}+n_{00})$. As there is no actual positive observations in Table 1(a), the true positive rate (TPR, i.e., power) is undefined for the simulation settings in Crawford et al. (2006). Therefore, the comparison of Type I error rates in Crawford et al. (2006) is not thoroughly discussed, as the study did not control for statistical power. Conversely, for the power analysis in Crawford and Garthwaite (2006), all single-subject observations are generated from a distribution that is different from the distribution of control group, see Table 1(b). In Table 1(b), $n_{11}$ is the number of true positive (TP) while $n_{01}$ is the number of false negative (FN). The true positive rate (TPR, i.e., power) is defined as $n_{11}/(n_{11}+n_{10})$. However, the Type I error is undefined. Again, Crawford and Garthwaite (2006) did not fully discuss the comparison of the power evaluation due to the absence of Type I error. Neither Crawford et al. (2006) nor Crawford and Garthwaite (2006) discussed that the ground true observations of a single-subject are well-mixed. For instance, if 50% of the single-subject observations are generated from underlying distribution of the control group data, the other 50% of observation comes from a different distribution, the Type I error and power of the $t$ score test would differ significantly. A direct evidence illustrated in Figure 2(c) shows that the powers (in median) of Crawford-Garthwaite (CG) method and BIGPAST under two-sided test are around 0.62 and 0.96 respectively. To address this issue, all the simulation settings (unless specified) of the single-subject in this paper are well-mixed, i.e., the single-subject observations are generated from a mixture of the null and alternative distribution. The well-mixed setting is more realistic and can provide a comprehensive evaluation of the performance of BIGPAST test under skewed Student’s $t$ assumption.
@@
Before delving into the rationale behind the assumption of the skewed Student’s $t$ distribution, it is essential to review the definitions of heavy tails and light tails within the context of the skewed Student’s $t$ distribution. Let $F(x)$ be the cumulative distribution function of random variable $X$. The right tail distribution is denoted as $\overline{F}(x)\coloneqq\Pr(X>x)$ while the left tail distribution is denoted as $\underline{F}(x)\coloneqq\Pr(X\leq x)$. $F$ is said to have a right heavy tail if $\lim_{x\to+\infty}\exp({tx})\overline{F}(x)=\infty$ for all $t>0$; have a right light tail if $\exp({tx})\overline{F}(x)=O(1)$ for large positive $x$ and some constant $t>0$. Similarly, $F$ is said to have a left heavy tail if $\lim_{x\to-\infty}\exp({tx})\underline{F}(x)=\infty$ for all $t>0$; have a left light tail if $\exp({tx})\underline{F}(x)=O(1)$ for small negative $x$ and some constant $t>0$. In other words, light tails decay to zero much faster than the exponential distribution, whereas heavy tails decay much slower, see Figure 1.
@@
Figure 1: The heavy and light tails of the standard skewed Student’s $t$ distribution for both positive and negative skewness parameters. The orange lines represent the standard Student’s $t$ density function, the green lines represent the curve of $\exp({-{\left\lvert x\right\rvert}})$. The vertical lines are quantiles: $t_{0.05},t_{0.95}$ (resp. $t_{0.05}^{*},t_{0.95}^{*}$) are 5% and 95% quantiles of the standard Student’s $t$ distribution (resp. the standard skewed Student’s $t$ distribution). In (a), $t_{0.05}=-2.353,t_{0.95}=2.353$, $t_{0.05}^{*}=-0.398,t_{0.95}^{*}=3.169$. In (b), $t_{0.05}=-2.015,t_{0.95}=2.015$, $t_{0.05}^{*}=-2.568,t_{0.95}^{*}=0.370$.
@@
It is well-known that the Student’s $t$ distribution is symmetric and has two heavy tails at both sides. The skewed Student’s $t$ distribution is a generalisation of the Student’s $t$ distribution (Azzalini and Capitanio, 2014). The skewness parameter $\alpha$ in Equation 2 changes the symmetry of Student’s $t$ distribution given the degree of freedom $\nu$. When $\alpha=0$ and $\nu$ is finite, the skewed Student’s $t$ distribution reduces to Student’s $t$ distribution which has both heavy tails; When $\alpha=0$ and $\nu=+\infty$, the skewed Student’s $t$ distribution reduces to normal distribution which has both lighter tails; When $|\alpha|=\infty$ and $\nu<+\infty$, the skewed Student’s $t$ distribution reduces to half Student’s $t$ distribution which can be tuned to allow for a heavy tail ($\nu\to 1$) or a light tail (large $\nu$). Therefore, for other settings of ($\alpha,\nu$), the skewed Student’s $t$ distribution may have one light tail and one heavy tail. The sign of $\alpha$ dictates whether the heavy tail is on the left or right side of the distribution, see Figure 1. In Figure 1(a), the skewness is positive, the standard skewed Student’s $t$ distribution has a right heavy tail and a left light tail. Conversely, in Figure 1(b), the skewness is negative, the standard skewed Student’s $t$ distribution has a right light tail and a left heavy tail, see the zoomed insets in Figures 1(a) and 1(b) respectively.
@@
In this paper, we consider three kinds of alternative hypothesis test: “less than” ($<$), “greater than” ($>$) and “two-sided”. For clarity, we refer to the alternative hypothesis with symbol ($<$) as the one-sided test (less), and the alternative hypothesis with symbol ($>$) as the one-sided test (greater). Under the assumption of symmetric distributions (e.g. Normal or Student’s $t$), one-sided test, either “less” or “greater”, share the same absolute critical value given the same significance level. However, under the assumption of skewed Student’s $t$ distribution, the absolute critical value of the one-sided test (less) is different from the absolute critical value of the one-sided test (greater) due to relationship between skewness and heavy tail. Hence, under the assumption of a skewed Student’s $t$ distribution, there are six combinations of hypothesis tests: the one-sided test (less), the one-sided test (greater), and the two-sided test, each with either a positive or negative skewness parameter.
@@
Now, let us look into the parameter settings in Crawford et al. (2006). They set $\alpha<0$ and only considered the one-sided test (less). Both the observations of control group and single-subject are generated from the same skewed Student’s $t$ distribution with the same skewness parameter. Furthermore, the observations of single-subject are all negative as described in Table 1(a). They used the $t$ statistic to test the abnormalities of the single-subject. Figure 1(b) describes a similar skewness parameter setting as in Crawford et al. (2006). Given the significance level 0.95, the critical value of the one-sided test (less) for $t$ test is $t_{0.05}=-2.015$ while the critical value of the one-sided test (less) for skewed $t$ test is $t_{0.05}^{*}=-2.568$. Apparently, $t_{0.05}>t^{*}_{0.05}$, which implies that without considering the randomness, if some single-subject observation $x^{*}$ is claimed as abnormal by $t_{0.05}^{*}$, then $x^{*}$ must be claimed as abnormal by $t_{0.05}$. Let’s define the model mis-specification error for one-sided test (less) as the probability $\Pr_{t}(t_{0.05}^{*}\leq x\leq t_{0.05})$ where $\Pr_{t}(\cdot)$ is the probability under the Student’s $t$ distribution. In Figure 1(b), the model mis-specification error for one-sided test (less) is $\Pr_{t}(-2.568\leq x\leq-2.015)\approx 0.0249$. However, if we consider the one-sided test (greater) under the same settings, the model mis-specification error for one-sided test (greater) is $\Pr_{t}(0.0370\leq x\leq 2.015)\approx 0.3131$. The model mis-specification error for the one-sided test (greater) is almost 12 times higher than that for the one-sided test (less). This means that the one-sided test (less) is more robust than the one-sided test (greater) when $\alpha<0$. Under the settings of $\alpha<0$, one-sided test (less) and not well-mixed observations of single-subject, Crawford et al. (2006) claimed that the $t$ score is robust in terms of Type I error even under skewed Student’s $t$ distribution. However, this assertion does not hold if the one-sided test (less) is replaced with the one-sided test (greater). For the power comparison, Crawford and Garthwaite (2006) make the similar statement as well. However, by using the similar discussion above, we can find that the robustness statement for the power only holds for the specific settings in their paper.
@@
Moreover, we investigate the model misspecification error between normal and skewed Student’s $t$ distributions as well in Figure 3. The regions shaded in light-red colour in Figures 3(a) and 3(b) represent the model misspecification errors that occur when the skewed Student’s $t$ distribution is assumed, but the $z$ score test is used. Apparently, the $z$ score will always have model misspecification error under the skewed Student’s $t$ distribution assumption. In fact, neither the $z$ score test, the $t$ score test, nor even the Crawford-Garthwaite Bayesian method (Crawford and Garthwaite, 2007) can eliminate the model misspecification error.
@@
The rationale for this paper’s assumption of the skewed Student’s $t$ distribution can be attributed to the fact that it can help eliminate model misspecification errors, provided that the underlying distribution is a special case of the skewed Student’s $t$ distribution. These special cases encompass the regular Student’s $t$ distribution ($\alpha=0$), skewed normal distribution ($\nu=\infty$), half normal distribution ($\alpha=\pm\infty,\nu=\infty$), and normal distribution ($\alpha=0,\nu=\infty$) as special cases. To show the elimination of model misspecification error, simulation results with normal assumption but using BIGPAST method can be found in Table 9.
@@
We can conclude that the interplay between skewness (whether positive or negative), the type of hypothesis (‘two-sided’, ‘less’, or ‘greater’), and the single-subject observations (whether well-mixed or not) significantly influences both the Type I error rate and the power of the $t$ score test under the assumption of skewed Student’s $t$ distribution. To investigate the influence, this paper introduces a Bayesian Inference General Procedures for A Single-subject Test (BIGPAST) framework under the skewed Student’s $t$ assumption to fully discuss the test performance in different settings (i.e., the combinations of the sign of $\alpha$, type of hypothesis and well-mixed single-subject observations). The BIGPAST framework is analogous to the Bayesian inference framework proposed by Crawford and Garthwaite (Crawford and Garthwaite, 2007), with the key distinction being our assumption of the skewed Student’s $t$ distribution. Additionally, we propose a Jefferys prior which is defined by explicit mathematical formulas. We evaluate BIGPAST’s performance against the existing approaches through simulation studies. The results demonstrate that BIGPAST is robust against departures from normality and surpasses the existing methods based on the assumption of a normal distribution in terms of Type I error, power and accuracy.
@@
The remainder of the paper is organised as follows. In Section 2, we introduce the Bayesian inference procedures for comparing a single-subject with a control group under the skewed Student’s $t$ distribution. In Section 3, we conduct a comprehensive set of simulation studies to evaluate the performance of the proposed BIGPAST method. We then apply the BIGPAST method to a MEG dataset comprising an individual with mild traumatic brain injury and a control group in Section 4. Finally, we summarise our findings and discuss future research directions in Section 5.

### v4

@@ Abstract
Abnormality detection in identifying a single-subject which deviates from the majority of a control group dataset is a fundamental problem. Typically, the control group is characterised using standard Normal statistics, and the detection of a single abnormal subject is in that context. However, in many situations, the control group cannot be described by Normal statistics, making standard statistical methods inappropriate. This paper presents a Bayesian Inference General Procedures for A Single-Subject Test (BIGPAST) designed to mitigate the effects of skewness under the assumption that the dataset of the control group comes from the skewed Student $t$ distribution. BIGPAST operates under the null hypothesis that the single-subject follows the same distribution as the control group. We assess BIGPAST’s performance against other methods through simulation studies. The results demonstrate that BIGPAST is robust against deviations from normality and outperforms the existing approaches in accuracy. BIGPAST can reduce model misspecification errors under the skewed Student $t$ assumption. We apply BIGPAST to a Magnetoencephalography (MEG) dataset consisting of an individual with mild traumatic brain injury and an age and gender-matched control group, demonstrating its effectiveness in detecting abnormalities in a single-subject.
Keywords— Bayesian inference, Skewed Student’s $t$ distribution, Single-subject test, Magnetoencephalography (MEG), Jeffery’s prior, and False discovery rate
@@ 1 Introduction
Abnormality detection is a fundamental task in many scientific fields, including medicine, psychology, and neuroscience (Scholz and Stephens, 1987; Crawford and Howell, 1998; Crawford and Garthwaite, 2007). One example of such abnormality detection lies in the detection of traumatic brain injury (TBI), which is a fundamental problem, with over 2.8 million individuals a year in America presenting at an Emergency room (Taylor et al., 2017). We exemplify the use of skewed Student $t$ statistics in the study of brain activity as a potential biomarker for the identification of a brain injury. We use magnetoencephalography (MEG) to measure brain activity and compare single individual to normal controls from age and gender-matched cohorts. Existing literature typically assumes that the dataset from the control group follows a Normal distribution. However, extensive practical evidence shows that the Normal distribution assumption is often violated. This violation is frequently observed in brain Magnetoencephalography (MEG) datasets. Therefore, the current approaches using Normal statistics could not be directly applied. Unlike the existing literature, we assume that the dataset from the control group belongs to a more general distributional family - the skewed Student $t$ distribution, which includes the Normal distribution and Student $t$ as special cases. The benefit of this skewed Student $t$ distributional assumption is that it could cover a variety of empirical behaviours and is thus more suitable for practical needs. This generalisation is essential since many brain Magnetoencephalography (MEG) datasets show a significant level of asymmetry in the observation distribution and hence are highly skewed. It is also the case that the high skewness not only appears in a MEG dataset but also in other fields, such as cybersecurity (Kim and Song, 2023), medicine (Lu and Huang, 2017), and neuroscience (Buzsáki and Mizuseki, 2014) to name only a few.
Under the assumption of the skewed Student $t$ distribution in this article, we describe the statistical issues involved in the comparison of a single individual with a large control group. The statistical problems arise from the interplay between the mechanism of data observation from a single-subject, the type of hypothesis (two-sided, less, greater), and the sign of the skewness parameter, which we will discuss one by one in the following paragraphs.
Under the assumption that the dataset observed from the control group follows a Normal distribution, several methods exist to compare a single-subject against a control group. Crawford and Howell (1998) proposed a $t$-score test defined as follows:
@@
where $x^{*}$ represents the observation of the single-subject, $\bar{x}$ denotes the average of $n$ observations from the control group, and $s$ stands for the standard deviation of the dataset from the control group. The test score $t$ follows a Student $t$ distribution with $n-1$ degrees of freedom. Let $\beta$ be the prespecified significance level, define the $p$ value as $\mathbb{P}(t>|t_{1-\beta/2}(n-1)|)$ where $t_{1-\beta/2}(n-1)$ is the $(1-\beta/2)$-th critical value of Student $t$ distribution with $n-1$ degrees of freedom. If the $p$ value is less than $\beta$, then one can claim that $x^{*}$ has an abnormality. Moreover, under the assumption of a skewed Student $t$ distribution, Crawford et al. (2006) examined the effects of departures from normality on testing for a deficit in a single-subject via using a Leptokurtic distribution with different skewness parameters. They asserted that the $t$ score (Crawford and Howell, 1998) outperforms the $z$ score in terms of Type I error, and the $t$ score shows robustness as good as the result under the Normal assumption. Additionally, Crawford and Garthwaite (2006) investigated the statistical power of testing for an abnormality in a single-subject using Monte Carlo simulations. They found that the power of $t$ score performs slightly worse than the Normal assumption results.
The simulation results for Type I error in Crawford et al. (2006) and power in Crawford and Garthwaite (2006) led to the assertion that the $t$ score Crawford and Howell (1998) is robust even when the underlying distribution is skewed Student $t$ distribution. However, the simulation settings for generating the samples for a single-subject in Crawford et al. (2006)Crawford and Garthwaite (2006) are not comprehensive to evaluate the performance of the $t$ score under skewed Student $t$ distribution. The samples of the single-subject may come from either the distribution of the control group or a different distribution. Let $c:d$ be the ratio of the samples from the control group to those from a different distribution. In Crawford et al. (2006), the ratio is $100:0$; and in Crawford and Garthwaite (2006), the ratio is $0:100$. Both are special cases for evaluating the Type I error and Power, respectively.
Table 1: The data generation mechanism of single-subject observations in Crawford et al. (2006)Crawford and Garthwaite (2006) respectively. ‘NA’ means that the observations in the cell are not available. ‘Positive’ means the single-subject is claimed as abnormal, while ‘Negative’ means the single-subject is claimed as normal.
@@
For $\mathbf{c:d=100:0}$, Crawford et al. (2006) generates all single-subject observations from the distribution of the control group to evaluate the Type I error only, see Table 1. In Table 1, $n_{01}$ is the number of false positives (FP) while $n_{00}$ is the number of true negatives (TN). The false positive rate (FPR, i.e., Type I error) is estimated as $n_{01}/(n_{01}+n_{00})$. As there are no actual positive observations in Table 1, the true positive rate (TPR, i.e., power) can not be estimated given the simulation settings in Crawford et al. (2006). Therefore, the comparison of Type I error rates in Crawford et al. (2006) is not thoroughly discussed, as the study did not control for statistical power. Conversely, for $\mathbf{c:d=0:100}$Crawford and Garthwaite (2006) generate all single-subject observations from a distribution that is different from the distribution of the control group to evaluate the power only, see Table 1. In Table 1, $n_{11}$ is the number of true positives (TP) while $n_{01}$ is the number of false negatives (FN). The true positive rate (TPR, i.e., power) is estimated as $n_{11}/(n_{11}+n_{10})$. However, the Type I error is unavailable based on the observations. Again, Crawford and Garthwaite (2006) did not fully discuss the comparison of the power evaluation due to the absence of Type I error. Neither Crawford et al. (2006) nor Crawford and Garthwaite (2006) discussed that the ground truth observations of a single-subject are well-mixed. For instance, if 50% of the single-subject observations are generated from the underlying distribution of the control group data and the other 50% of observations come from a different distribution, the Type I error and power of the $t$ score test would differ significantly. Direct evidence illustrated in Figure 2(c) shows that the median of powers of the Crawford-Garthwaite (CG) method and BIGPAST under two-sided test are around 0.62 and 0.96, respectively. To address this issue, all the simulation settings (unless specified) of the single-subject in this paper are well-mixed, i.e., the single-subject observations are generated from a mixture of the null and alternative distribution. The well-mixed setting is more realistic and can provide a comprehensive evaluation of the performance of BIGPAST test under skewed Student $t$ assumption.
@@
Before delving into the rationale behind the assumption of the skewed Student $t$ distribution, it is essential to review the definitions of heavy tails and light tails within the context of the skewed Student $t$ distribution. Let $F(x)$ be the cumulative distribution function of a random variable $X$. The right tail distribution is denoted as $\overline{F}(x)\coloneqq\Pr(X>x)$ while the left tail distribution is denoted as $\underline{F}(x)\coloneqq\Pr(X\leq x)$. $F$ is said to have a right heavy tail if $\lim_{x\to+\infty}\exp({tx})\overline{F}(x)=\infty$ for all $t>0$, and have a right light tail if $\exp({tx})\overline{F}(x)=O(1)$ for large positive $x$ and some constant $t>0$. Similarly, $F$ is said to have a left heavy tail if $\lim_{x\to-\infty}\exp({tx})\underline{F}(x)=\infty$ for all $t>0$, and have a left light tail if $\exp({tx})\underline{F}(x)=O(1)$ for small negative $x$ and some constant $t>0$. In other words, light tails decay to zero much faster than the exponential distribution, whereas heavy tails decay much slower, see Figure 1.
@@
Figure 1: The heavy and light tails of the standard skewed Student $t$ distribution for both positive and negative skewness parameters. The orange lines represent the standard Student $t$ density function, the green lines represent the curve of $\exp({-{\left\lvert x\right\rvert}})$. The vertical lines are quantiles: $t_{0.05},t_{0.95}$ (resp. $t_{0.05}^{*},t_{0.95}^{*}$) are 5% and 95% quantiles of the standard Student $t$ distribution (resp. the standard skewed Student $t$ distribution). In (a), $t_{0.05}=-2.353,t_{0.95}=2.353$, $t_{0.05}^{*}=-0.398,t_{0.95}^{*}=3.169$. In (b), $t_{0.05}=-2.015,t_{0.95}=2.015$, $t_{0.05}^{*}=-2.568,t_{0.95}^{*}=0.370$.
It is well-known that the Student $t$ distribution is symmetric and has two heavy tails. The skewed Student $t$ distribution is a generalisation of the Student $t$ distribution (Azzalini and Capitanio, 2014). Let $f(z|\alpha,\nu)$ be the skewed Student $t$ density function (Azzalini and Capitanio, 2003), defined as follows:
where $r(z,\nu)=\sqrt{(\nu+1)/(\nu+z^{2})}$ and $w=\alpha zr(z,\nu)$, $t(\cdot|\nu)$ represents the density function of the $t$ distribution with degrees of freedom $\nu$, and $T(\cdot|\nu+1)$ represents the cumulative density function of the $t$ distribution with degrees of freedom $\nu+1$. Let $\xi$ and $\omega$ be the location and scale parameters, with the substitution $z=(x-\xi)/\omega$ in Equation 1, then the skewed Student $t$ density function is given by
@@
The skewness parameter $\alpha$ changes the symmetry of Student $t$ distribution given the degrees of freedom $\nu$. When $\alpha=0$, $\nu$ is finite, the skewed Student $t$ distribution reduces to Student $t$ distribution which has heavy tails. When $\alpha=0$ and $\nu\to+\infty$, the skewed Student $t$ distribution reduces to the Normal distribution which has lighter tails; When $|\alpha|\to\infty$ and $\nu<+\infty$, the skewed Student $t$ distribution reduces to the half Student $t$ distribution which can be tuned to allow for a heavy tail ($\nu\to 1$) or a light tail (large $\nu$). Therefore, for other settings of ($\alpha,\nu$), the skewed Student $t$ distribution may have one light tail and one heavy tail. The sign of $\alpha$ dictates whether the heavy tail is on the left or right side of the distribution, see Figure 1. In Figure 1(a), the skewness is positive, and the standard skewed Student $t$ distribution has a right heavy tail and a left light tail. Conversely, in Figure 1(b), the skewness is negative, the standard skewed Student $t$ distribution has a right light tail and a left heavy tail, see the zoomed insets in Figures 1(a) and 1(b) respectively.
@@
In this paper, we consider three kinds of alternative hypotheses: “less than” ($<$), “greater than” ($>$) and “two-sided”. For clarity, we refer to the alternative hypothesis with symbol ($<$) as the one-sided test (less), and the alternative hypothesis with symbol ($>$) as the one-sided test (greater). Under the assumption of symmetric distributions (e.g., Normal or Student $t$), one-sided tests, either “less” or “greater”, share the same absolute critical value given the same significance level. However, under the assumption of the skewed Student $t$ distribution, the absolute critical value of the one-sided test (less) is different from the absolute critical value of the one-sided test (greater) due to the relationship between skewness and heavy tails. Hence, under the assumption of a skewed Student $t$ distribution, there are six combinations of hypothesis tests: the one-sided test (less), the one-sided test (greater), and the two-sided test, each with either a positive or negative skewness parameter.
@@
Let us look into the parameter settings in Crawford et al. (2006). They set $\alpha<0$ and only considered the one-sided test (less). The observations of the control group and single-subject are generated from the same skewed Student $t$ distribution with the same skewness parameter. Furthermore, the observations of the single-subject are all negative, as described in Table 1. They used the $t$ statistic to test the abnormalities of the single-subject. Figure 1(b) describes a similar skewness parameter setting as in Crawford et al. (2006). Given the significance level 0.05, the critical value of the one-sided test (less) for the $t$ test is $t_{0.05}=-2.015$ while the critical value of the one-sided test (less) for the skewed $t$ test is $t_{0.05}^{*}=-2.568$. Apparently, $t_{0.05}>t^{*}_{0.05}$, which implies that if some single-subject observation $x^{*}$ is claimed as abnormal by $t_{0.05}^{*}$, then $x^{*}$ must be claimed as abnormal by $t_{0.05}$.
Let’s define the model misspecification error for the one-sided test (less) as the probability $\Pr_{t}(t_{0.05}^{*}\leq x\leq t_{0.05})$ where $\Pr_{t}(\cdot)$ is the probability under the Student $t$ distribution. In Figure 1(b), the model misspecification error for the one-sided test (less) is $\Pr_{t}(-2.568\leq x\leq-2.015)\approx 0.0249$. However, if we consider the one-sided test (greater) under the same settings, the model misspecification error for the one-sided test (greater) is $\Pr_{t}(0.0370\leq x\leq 2.015)\approx 0.3131$. The model misspecification error for the one-sided test (greater) is almost 12 times higher than that for the one-sided test (less). This means that the one-sided test (less) is more robust than the one-sided test (greater) when $\alpha<0$. Under the settings of $\alpha<0$, one-sided test (less) and not well-mixed observations of single-subject, Crawford et al. (2006) claimed that the $t$ score is robust in terms of Type I error even under skewed Student $t$ distribution. However, this assertion does not hold if the one-sided test (less) is replaced with the one-sided test (greater). For the power comparison, Crawford and Garthwaite (2006) make a similar statement. However, by using a similar discussion to the above, we can find that the robustness statement for the power only holds for the specific settings in their paper.
@@
Moreover, we also investigate the model misspecification error between Normal and skewed Student $t$ distributions in Figure 4. The $z$ score will always have model misspecification error under the skewed Student $t$ distribution assumption. Neither the $z$ score test, the $t$ score test, nor even the Crawford-Garthwaite Bayesian method (Crawford and Garthwaite, 2007) can eliminate the model misspecification error.
@@
The rationale for this paper’s assumption of the skewed Student $t$ distribution for the underlying proposal can be attributed to the fact that it can help eliminate model misspecification errors, provided that the underlying distribution is a special case of the skewed Student $t$ distribution. These special cases encompass the regular Student $t$ distribution ($\alpha=0$), skewed Normal distribution ($\nu\to\infty$), half Normal distribution ($\alpha\to\pm\infty,\nu\to\infty$), and Normal distribution ($\alpha=0,\nu\to\infty$) as special cases. To show the elimination of model misspecification error, simulation results with a Normal assumption but using BIGPAST method can be found in Table 10.
@@
We can conclude that the interplay between skewness (whether positive or negative), the type of hypothesis (‘two-sided’, ‘less’, or ‘greater’), and the single-subject observations (whether well-mixed or not) significantly influences both the Type I error rate and the power of the $t$ score test under the assumption of the skewed Student $t$ distribution. The key contribution is that this paper introduces a Bayesian Inference General Procedures for A Single-subject Test (BIGPAST) framework, which employs the nested sampling technique under the skewed Student $t$ assumption to fully discuss the test performance in different settings (i.e., the combinations of the sign of $\alpha$, type of hypothesis and well-mixed single-subject observations).
Additionally, another contribution of this paper lies in the fact that we propose a Jeffreys prior, which is defined by explicit mathematical formulas. We evaluate BIGPAST’s performance against the existing approaches through simulation studies. The results demonstrate that BIGPAST is robust against departures from normality and surpasses the existing methods based on the assumption of a Normal distribution in terms of Type I error, power and accuracy.
@@
The remainder of the paper is organised as follows. In Section 2, we introduce the Bayesian inference procedures for comparing a single-subject with a control group under the skewed Student $t$ distribution. In Section 3, we conduct a comprehensive set of simulation studies to evaluate the performance of the proposed BIGPAST method. In Section 4, we apply the BIGPAST method to a MEG dataset comprising an individual with mild traumatic brain injury and a control group. Finally, we summarise our findings and discuss future research directions in Section 5.